OpenStudy (anonymous):

Determine the number of solutions of 2x2 + 3x = 0.

OpenStudy (shayaan_mustafa):

Hi iceyweheniplease7 :) How are you?

OpenStudy (shayaan_mustafa):

Your equation is look like this, \[\Large 2x^2+3x=0\]

OpenStudy (hba):

4+3x=0 3x=-4 x=-4/3

OpenStudy (shayaan_mustafa):

It is a two degree equation so it has two solutions.

OpenStudy (anonymous):

IM GREAT :) and how r u?:)

OpenStudy (shayaan_mustafa):

The highest degree is 2

OpenStudy (shayaan_mustafa):

Fine. Thank you :)

OpenStudy (anonymous):

2x^2+3x=2x(x+3/2)----->x=0 & x=-3/2

OpenStudy (anonymous):

the degree of the equation means nothing as to whether or not it has any solution. anyways...yours has one solution. x=-3/2

OpenStudy (anonymous):

the degree of the equation does tell yo the max number of POSSIBLE zeros. that's about it...

OpenStudy (shayaan_mustafa):

@hba you are wrong my friend.

OpenStudy (hba):

@Shayaan_Mustafa i missunderstood the question :D

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

lol...forgot the trivial solution. Yeah x=0 works too :)

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

thanks for the reminder regarding division by zero :)